# PV-SSD: A Multi-Modal Point Cloud 3D Object Detector based on Projection Features and Voxel Features

### v5

@@ II-B Projection-based 3D object detection method
The projection-based methods typically project unordered point clouds into different views, transform them into regular pseudo-image representations, and complete subsequent feature extraction and detection tasks using 2D convolution. Based on the feature encoding method, projection-based methods can be classified into two categories. The first category, such as MV3D [34], AVOD [35], Complex YOLO [21], and YOLO3D [24], project point clouds into height, density, and intensity maps for feature extraction. MV3D and AVOD employ multi-view fusion, projecting point clouds into bird’s-eye and perspective views. Complex-YOLO and YOLO3D use single-view projection, projecting point clouds into a bird’s-eye view. The second category converts point clouds into pillar structures. PointPillars [23] first proposed this encoding method, which converts point clouds into a top-down pillar structure and utilizes this structure to perform feature encoding. H3D R-CNN [29], following the same feature encoding method as PointPillars, reconstructs feature points from bird’s-eye and perspective views into a 3D feature representation. This structure fully utilizes the complementary nature of bird’s-eye and perspective views, achieving better detection accuracy. However, projecting point clouds into a specific view often leads to excessive compression of point cloud information, resulting in a significant loss of feature information.
@@ II-C Voxel-based 3D object detection method
To handle unstructured point clouds, VoxelNet [25] first proposed a method that converts unordered point clouds into regular voxels and then extracts features using 3D convolutional neural networks. However, in most cases in autonomous driving scenarios, the majority of the generated voxels are filled with zeros. Extracting features using 3D convolution would result in a significant amount of ineffective computation. To address this issue, SECOND [26] introduced the sparse convolution, which greatly improves computational efficiency. Voxel R-CNN [36] extends SECOND into a two-stage method and proposes Voxel RoI pooling to aggregate 3D features for refining region proposals. This two-stage structure allows Voxel R-CNN to achieve good detection accuracy even in coarse-grained voxels. In subsequent work, it was found that in autonomous driving scenarios, it is possible to obtain not only the annotated bounding boxes and the foreground points they contain, but also the spatial distribution information of each foreground. Leveraging this information can greatly enhance the accuracy of boundary box prediction. In response to this finding, PartA [37] proposed a new two-stage architecture. It incorporates the prediction of the internal spatial distribution information of foreground objects in the first stage, thereby achieving high-quality 3D bounding box detection. Subsequently, CIA-SSD [38] and SA-SSD [39] address the problem of misalignment between bounding boxes and classification confidence. CIA-SSD uses an IoU-aware confidence correction module and a multi-task head module to improve the imbalance between classification confidence and regression confidence in 3D detection tasks. SA-SSD proposes an auxiliary network that can be turned off at any time to enhance the network’s sensitivity to boundary information. In autonomous driving scenarios, the distribution of point clouds is sparse and uneven. This distribution poses challenges for point cloud sampling. To address this issue, SARPNET [40] proposes a new low-level feature encoder and utilizes a shape attention mechanism to learn the 3D shape priors of objects. Voxel-based methods show outstanding performance in terms of accuracy and speed due to the superiority of the convolution structure, but the choice of voxel size is still sensitive. To address this problem, SMS-Net [31] proposes a sparse multi-scale voxel feature aggregation network, which divides the raw point cloud into voxels of different sizes to perceive 3D features in different perceptual fields. Although voxel-based methods can preserve more point cloud information compared to projection, voxel-based methods often suffer from the loss of fine-grained feature information due to downsampling.
@@ IV-C Compared with Others
As can be seen from Table \@slowromancapi@, our method has 3% lower 3D AP in the ‘Car’ class compared to the H3D R-CNN[29], which is also based on the projection method. After our analysis, the reasons are: (1) The H3D R-CNN uses a dynamic voxelization feature coding method [60], establishing a complete mapping relationship between points and voxels. In contrast, we use a hard voxelization approach [25], which leads to uncertain voxel filling and information loss due to random sampling and discarding operations. (2) The projected features in the H3D R-CNN retain higher dimensional feature information through the MLP layer. In contrast, we use the same feature coding method as Complex-YOLO [21], which over-compresses the features in the bird’s-eye view, resulting in information loss. (3) H3D R-CNN is a two-stage 3D detector that fully uses the prior knowledge of the Region Proposal Network (RPN) module. In contrast, our method is a one-stage 3D detector, and the detection accuracy depends heavily on how well the features are extracted. However, we found that our method has good accuracy for objects with fewer reflection points like cyclists. We analyze the main reasons for this as: (1) However, the hard voxelization we use causes some information loss, we use a weight-based feature point sampling method in the Re-voxelization layer. For objects with fewer reflective points, such as cyclists, we can prioritize the feature points that are more beneficial to the detection task. (2) PWF-Net can filter out the point feature information more beneficial to the detection task. (3) For objects with few reflection points and small sizes, like the ‘Cyclist’ class, the voxel description in 3D space is more responsive to its features than the 2D perspective view. In the MSSFA module, we complemented the BEV-map features with Voxel-BEV features at the same grid points. In the subsequent ablation experiments, we can also see that the Re-voxelization layer, PWF-Net and MSP-Fusion module have a good improvement of the detection accuracy for the ‘Cyclist’ class.
@@ V Conclusion
Furthermore, when extracting voxel features using convolutional neural networks, there is a loss of fine-grained crucial information due to downsampling. This information loss has a significant impact on objects with fewer reflection points, such as cyclists. We propose a sampling method based on feature point weights. Subsequent experiments have also found that although there is a gap in detection accuracy for cars compared to the latest work H3D R-CNN, our method has a good competitiveness in detecting cyclists. In addition, in the ablation experiments, we compare the detection results using random sampling and feature point weight-based sampling. It is found that the method based on feature point weight sampling achieved a 2%-3% improvement.

### v6

@@ II-B Projection-based 3D object detection method
The projection-based methods typically project unordered point clouds into different views, transform them into regular pseudo-image representations, and complete subsequent feature extraction and detection tasks using 2D convolution. Based on the feature encoding method, projection-based methods can be classified into two categories. The first category, such as MV3D [34], AVOD [35], Complex YOLO [21], and YOLO3D [24], project point clouds into height, density, and intensity maps for feature extraction. MV3D and AVOD employ multi-view fusion, projecting point clouds into bird’s-eye and perspective views. Complex-YOLO and YOLO3D use single-view projection, projecting point clouds into a bird’s-eye view. The second category converts point clouds into pillar structures. PointPillars [23] first proposed this encoding method, which converts point clouds into a top-down pillar structure and utilizes this structure to perform feature encoding. H23D R-CNN [29], following the same feature encoding method as PointPillars, reconstructs feature points from bird’s-eye and perspective views into a 3D feature representation. This structure fully utilizes the complementary nature of bird’s-eye and perspective views, achieving better detection accuracy. However, projecting point clouds into a specific view often leads to excessive compression of point cloud information, resulting in a significant loss of feature information.
@@ II-C Voxel-based 3D object detection method
To handle unstructured point clouds, VoxelNet [25] first proposed a method that converts unordered point clouds into regular voxels and then extracts features using 3D convolutional neural networks. However, in most cases in autonomous driving scenarios, the majority of the generated voxels are filled with zeros. Extracting features using 3D convolution would result in a significant amount of ineffective computation. To address this issue, SECOND [26] introduced the sparse convolution, which greatly improves computational efficiency. Voxel R-CNN [36] extends SECOND into a two-stage method and proposes Voxel RoI pooling to aggregate 3D features for refining region proposals. This two-stage structure allows Voxel R-CNN to achieve good detection accuracy even in coarse-grained voxels. In subsequent work, it was found that in autonomous driving scenarios, it is possible to obtain not only the annotated bounding boxes and the foreground points they contain, but also the spatial distribution information of each foreground. Leveraging this information can greatly enhance the accuracy of boundary box prediction. In response to this finding, PartA2 [37] proposed a new two-stage architecture. It incorporates the prediction of the internal spatial distribution information of foreground objects in the first stage, thereby achieving high-quality 3D bounding box detection. Subsequently, CIA-SSD [38] and SA-SSD [39] address the problem of misalignment between bounding boxes and classification confidence. CIA-SSD uses an IoU-aware confidence correction module and a multi-task head module to improve the imbalance between classification confidence and regression confidence in 3D detection tasks. SA-SSD proposes an auxiliary network that can be turned off at any time to enhance the network’s sensitivity to boundary information. In autonomous driving scenarios, the distribution of point clouds is sparse and uneven. This distribution poses challenges for point cloud sampling. To address this issue, SARPNET [40] proposes a new low-level feature encoder and utilizes a shape attention mechanism to learn the 3D shape priors of objects. Voxel-based methods show outstanding performance in terms of accuracy and speed due to the superiority of the convolution structure, but the choice of voxel size is still sensitive. To address this problem, SMS-Net [31] proposes a sparse multi-scale voxel feature aggregation network, which divides the raw point cloud into voxels of different sizes to perceive 3D features in different perceptual fields. Although voxel-based methods can preserve more point cloud information compared to projection, voxel-based methods often suffer from the loss of fine-grained feature information due to downsampling.
@@ IV-C Compared with Others
As can be seen from Table \@slowromancapi@, our method has 3% lower 3D AP in the ‘Car’ class compared to the H23D R-CNN[29], which is also based on the projection method. After our analysis, the reasons are: (1) The H23D R-CNN uses a dynamic voxelization feature coding method [60], establishing a complete mapping relationship between points and voxels. In contrast, we use a hard voxelization approach [25], which leads to uncertain voxel filling and information loss due to random sampling and discarding operations. (2) The projected features in the H23D R-CNN retain higher dimensional feature information through the MLP layer. In contrast, we use the same feature coding method as Complex-YOLO [21], which over-compresses the features in the bird’s-eye view, resulting in information loss. (3) H23D R-CNN is a two-stage 3D detector that fully uses the prior knowledge of the Region Proposal Network (RPN) module. In contrast, our method is a one-stage 3D detector, and the detection accuracy depends heavily on how well the features are extracted. However, we found that our method has good accuracy for objects with fewer reflection points like cyclists. We analyze the main reasons for this as: (1) However, the hard voxelization we use causes some information loss, we use a weight-based feature point sampling method in the Re-voxelization layer. For objects with fewer reflective points, such as cyclists, we can prioritize the feature points that are more beneficial to the detection task. (2) PWF-Net can filter out the point feature information more beneficial to the detection task. (3) For objects with few reflection points and small sizes, like the ‘Cyclist’ class, the voxel description in 3D space is more responsive to its features than the 2D perspective view. In the MSSFA module, we complemented the BEV-map features with Voxel-BEV features at the same grid points. In the subsequent ablation experiments, we can also see that the Re-voxelization layer, PWF-Net and MSP-Fusion module have a good improvement of the detection accuracy for the ‘Cyclist’ class.
@@ V Conclusion
Furthermore, when extracting voxel features using convolutional neural networks, there is a loss of fine-grained crucial information due to downsampling. This information loss has a significant impact on objects with fewer reflection points, such as cyclists. We propose a sampling method based on feature point weights. Subsequent experiments have also found that although there is a gap in detection accuracy for cars compared to the latest work H23D R-CNN, our method has a good competitiveness in detecting cyclists. In addition, in the ablation experiments, we compare the detection results using random sampling and feature point weight-based sampling. It is found that the method based on feature point weight sampling achieved a 2%-3% improvement.